A recent survey examined the use of social media platforms. Suppose the survey found that there is a 0.63 probability that a randomly selected person will use Facebook and a 0.26 probability that a randomly selected person will use LinkedIn. In addition, suppose there is a 0.24 probability that a randomly selected person will use both Facebook and LinkedIn.
(a)
What is the probability that a randomly selected person will use Facebook or LinkedIn?
(b)
What is the probability that a randomly selected person will not use either social media platform?

Answer by ikleyn(53937)

.

(a)  P(F or L) = P(F) + P(L) - P(F and L) = 0.63 + 0.26 - 0.24 = 0.65.   ANSWER



(b)  Probability (b) is the COMPLEMENT to the probability (a)  1-0.65 = 0.35.    ANSWER